# LaMET’s Asymptotic Extrapolation vs. Inverse Problem

### v1

@@ Abstract
Large-Momentum Effective Theory (LaMET) is a physics-guided systematic expansion to calculate light-cone parton distributions, including collinear (PDFs) and transverse-momentum-dependent ones, at any fixed momentum fraction $x$ within a range of $[x_{\rm min},x_{\rm max}]$. It theoretically solves the ill-posed inverse problem that afflicts other theoretical approaches to collinear PDFs, such as short-distance factorizations. Recently, arXiv:2504.17706 [1] raised practical concerns about whether current or even future lattice data will have sufficient precision in the sub-asymptotic correlation region to support an error-controlled extrapolation—and if not, whether it becomes an inverse problem where the relevant uncertainties cannot be properly quantified. While we agree that not all current lattice data have the desired precision to qualify for an asymptotic extrapolation, some calculations do, and more are expected in the future. We comment on the analysis and results in Ref. [1] and argue that a physics-based systematic extrapolation still provides the most reliable error estimates, even when the data quality is not ideal. In contrast, re-framing the long-distance asymptotic extrapolation as a data-driven-only inverse problem with ad hoc mathematical conditioning could lead to unnecessarily conservative errors.
@@ I The Problem
Over the past decade, there has been significant progress in the lattice QCD calculation of parton physics, including parton distribution functions (PDFs), generalized parton distributions (GPDs), and transverse-momentum-dependent distributions (TMDs), etc. During this period, quite a few methods [2, 3, 4, 5, 6, 7, 8, 9, 10] have been developed to complement the traditional approaches of computing the lowest few moments of the parton distributions, aiming at either higher moments or the $x$-dependence of parton distributions. Among them, Large-Momentum Effective Theory (LaMET) [6, 11, 12], also known as the quasi-PDF method in the calculation of collinear PDFs, provides the most general framework for computing PDFs, GPDs, and TMDs, as well as light-cone wave functions. It has so far been the most widely adopted approach, achieving significant successes across all frontiers of parton physics [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 12, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142].
@@
LaMET is a physics-guided effective field theory (EFT) expansion of light-cone parton distributions, including the PDFs, GPDs, and TMDs, at any fixed momentum fraction $x$ within a moderate range of $[x_{\rm min},x_{\rm max}]$ [143], in terms of the quark and gluon momentum distributions in a fast moving hadron, aka the quasi distributions [6]. As an important advantage, it bypasses the inverse problem of reconstructing the PDF $x$-dependence from correlations of limited range, which afflicts other theoretical approaches such as the short-distance factorization (SDF), including the pseudo-PDF [7, 144] and “good lattice cross section” [8]. The SDF approach has also gained notable traction and led to substantial progress in the extracting the lowest Mellin moments and phenomenological fitting of collinear PDFs and GPDs [145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 67, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 144, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 173, 191, 192, 193], despite challenges in controlling the model uncertainty and its limited applicability to a broader range of parton physics.
@@
Figure 1: An illustration of the difference between LaMET and SDF: LaMET in principle utilizes data at all $z$ to sharply resolve the momentum of a parton, though in practice they are usually truncated at $z_{L}\sim 1.0$ fm due to noisy data and extrapolated using physics-motivated asymptotic forms to $z=\infty$ [84]. Meanwhile, the SDF can only use data up to about $z_{\rm max}\sim 0.2-0.3$ fm (shaded blue region) to fit a model of the PDF, which constitutes an inverse problem. Plotted are the hybrid-scheme nucleon transversity quasi-PDF matrix elements computed in Ref. [89]. The shaded-red ellipse approximately indicates the sub-asymptotic region where the exponential decay starts to dominate. The concern raised in  [1] is whether the data precision in this region allows for a reliable error estimate of the extrapolation; otherwise, it must be treated as an inverse problem.
@@
where $|P\rangle$ is a plane-wave hadron state with momentum $P^{\mu}=(P^{t},0,0,P^{z})$, and $W(z,0)$ with $z^{\mu}=(0,0,0,z)$ is a spatial Wilson line that guarantees the gauge invariance of the correlator. In SDF, the data contain leading-twist information up to about $z=0.2-0.3$ fm, beyond which the framework breaks down. For instance, the average size of the instanton, a non-perturbative gluon configuration, is about 0.3 fm [194, 195]. At finite momentum $P^{z}$, only a limited range of leading-twist correlations can be accessed, which is insufficient for a reliable forward determination of the PDFs. Instead, this segment of correlations can be used for phenomenological fittings of the PDFs. In contrast, LaMET legitimately utilizes matrix elements at all $z$ to compute the quasi-PDF, thereby providing a sharply-defined momentum fraction $x$ for quarks and gluons. This explains why LaMET can calculate the PDFs at a fixed momentum fraction $x$ through a systematic expansion. This distinction between LaMET and SDF has been extensively discussed in the literature [12, 84, 196, 197, 143], where it is argued that, given the same lattice data at a large $P^{z}$, LaMET offers the most accurate determination of a PDF within a moderate range of $x$.
Recently, Ref. [1] raised practical concerns about whether lattice data in the sub-asymptotic region, i.e., $0.7\leavevmode\nobreak\ \text{fm}\lesssim z\lesssim 1.0\leavevmode\nobreak\ \text{fm}$ presented there, are precise enough for a controlled physical extrapolation to larger $z$. If not, this could adversely impact the Fourier transform (FT) used to obtain quasi-PDFs for the LaMET expansion, rendering the standard error estimation in the literature possibly unreliable. Various data-driven inverse-problem (IP) methods [198, 199, 200, 201, 202] were proposed to treat the FT, yielding large uncertainty of the constructed PDF, dubbed as “inverse problem in the LaMET framework”.
In this paper, we directly respond to these concerns. In Sec. II, we first briefly review the fundamental difference in the predictive capabilities of LaMET and SDF, based on the extensive discussions in Refs. [84, 12, 196, 197, 143], emphasizing LaMET as a forward-problem (FP) formalism of calculating, not fitting, parton distributions. Then in Sec. III we point out that the alleged IP in LaMET application to lattice data is a standard asymptotic extrapolation problem in lattice QCD, which can be systematically treated with reliable error estimates. In Sec. IV, we examine in detail the criticisms and data analysis in Ref. [1], and conclude that the observed FT errors can be bounded by a theoretical estimate based on asymptotic extrapolation, and the data-driven IP methods used are less reliable as they do not satisfy the physical constraints. We conclude in Sec. V.
II LaMET As a Forward Problem Formalism
@@
As presented in Eq. (3), the alleged inverse problem in Ref. [1] is essentially a matter of convergence and precision of the numerical Fourier integral, which are influenced by the discretization and and the extrapolation in $z$-space. The smoothness of correlation functions in a bounded hadron ensures that the discretization effects are small, while their asymptotic behavior at large distances controls the extrapolation uncertainty. Therefore, according to Eqs. (II) and (3), LaMET is a forward problem formalism in lattice applications, which is systematically improvable without an adverse IP.
It is important to emphasize that in this formulation, the shapes and mathematical properties of the PDFs are not assumed a priori, in contrast to the IP formulation for SDF [1]. In fact, PDFs are calculated one $x$ at a time, and the $x$-space functions come out as a prediction. Since no connections with partons are made before the large-momentum expansion, and the terms in SDF for the coordinate-space correlations—such as “higher-twist contamination at large $z$” [1]—are irrelevant. Moreover, unlike relevant statements in Ref. [1], the asymptotic behavior of the correlation functions will constrain any sharp structures in the PDFs—associated with long-range physics—to be fluctuations no larger than the extrapolation error. Besides, these fluctuations are also closely tied to the smallest value of $x$ that LaMET can probe, which is limited by the largest momentum $P^{z}$. As a result, any sharp fluctuations in the PDF are already bounded within the FT uncertainty and power corrections. Ultimately, LaMET’s ability to resolve such fine features will improve with higher precision and access to larger momenta.
@@
Here ${\cal C}$ is the matching kernel, and the power corrections are suppressed by the short distance $z$. After taking into account higher-order perturbative corrections [206, 76, 207], the SDF is found to be reliable up to $z=0.2-0.3$ fm, which corresponds to the strong coupling constant $\alpha_{s}(1/z)\sim 1$ and the breakdown of perturbation theory. As a result, the largest $\lambda$ that can be reached in contemporary lattice simulations is limited to $\lambda_{\rm max}\sim 3.0-4.0$. Therefore, the $x$-dependence of PDFs cannot be calculated directly and must instead be fitted through parametrizations and Bayesian or neural-network approaches with assorted numerical and mathematical assumptions [144, 173, 156, 188].
@@
To facilitate the LaMET analysis, one needs to have coordinate-space correlation functions in a large-momentum $P^{z}$ hadron at all distances $z\in[0,\infty]$, as information at negative $z$ can be obtained from discrete symmetries. It is well known that lattice QCD cannot produce quality data for $z\gg 1$ fm due to the exponentially worsening signal-to-noise ratio (SNR). Therefore, we cannot have all the data that LaMET requires. Fortunately, physical principles offer important guidance. Even without doing a very challenging calculation, QCD predicts that the correlation function must effectively vanish at sufficiently large $z$ due to color confinement. In fact, we know a lot better. According to dispersion theory, the correlation function decays exponentially at large $z$ with the decay length controlled by the residual mass of a heavy-light meson [65]. Moreover, this mass is an intrinsic property of the correlation function in the sense that it is independent of the external hadron momentum. Through more detailed analysis [208], one can derive subleading behaviors of the correlation function as well. Armed with the known rigid physics constraints, the lattice data can be reliably extrapolated to larger $z$ in the asymptotic region.
@@ III.1 Asymptotic Extrapolation
Asymptotic expansion is essentially a generalization of the Taylor expansion in small parameters, which is widely used in theoretical physics, such as effective field theories. It is one of the most frequently used techniques in lattice QCD analyses, where the physical conditions are imposed as a prior. The precision of calculations is often assessed based on this procedure.
@@
Other examples include continuum and chiral extrapolations [212, 213, 214], where subleading effects from finite lattice spacing and unphysical pion masses are derived using effective field theory and incorporated into fits to lattice data. In the same spirit, since dispersion relation analysis is a rigorous and model-independent method in theoretical physics, it supports the use of asymptotic extrapolation for spatial correlation functions as a standard and justified procedure.

### v2

@@ Abstract
Large-Momentum Effective Theory (LaMET) is a physics-guided systematic expansion to calculate light-cone parton distributions, including collinear (PDFs) and transverse-momentum-dependent ones, at any fixed momentum fraction $x$ within a range of $x_{\rm min},x_{\rm max}]$. It theoretically solves the ill-posed inverse problem that afflicts other theoretical approaches to collinear PDFs, such as short-distance factorizations. Recently, arXiv:2504.17706 [1] raised practical concerns about whether current or even future lattice data will have sufficient precision in the sub-asymptotic correlation region to support an error-controlled extrapolation—and if not, whether it becomes an inverse problem where the relevant uncertainties cannot be properly quantified. While we agree that not all current lattice data have the desired precision to qualify for an asymptotic extrapolation, some calculations do, and more are expected in the future. We comment on the analysis and results in Ref. [1] and argue that a physics-based systematic extrapolation still provides the most reliable error estimates, even when the data quality is not ideal. In contrast, re-framing the long-distance asymptotic extrapolation as a data-driven-only inverse problem with ad hoc mathematical conditioning could lead to unnecessarily conservative errors.
@@ I The Problem
Over the past decade, there has been significant progress in the lattice QCD calculation of parton physics, including parton distribution functions (PDFs), generalized parton distributions (GPDs), and transverse-momentum-dependent distributions (TMDs), etc. During this period, quite a few methods [2, 3, 4, 5, 6, 7, 8, 9, 10] have been developed to complement the traditional approaches of computing the lowest few moments of the parton distributions, aiming at either higher moments or the $x$-dependence of parton distributions. Among them, Large-Momentum Effective Theory (LaMET) [6, 11, 12], also known as the quasi-PDF method in the calculation of collinear PDFs, provides the most general framework for computing PDFs, GPDs, and TMDs, as well as light-cone wave functions. It has so far been the most widely adopted approach, achieving significant successes across all frontiers of parton physics [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 12, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171].
@@
LaMET is a physics-guided effective field theory (EFT) expansion of light-cone parton distributions, including the PDFs, GPDs, and TMDs, at any fixed momentum fraction $x$ within a moderate range of $x_{\rm min},x_{\rm max}] [$172], in terms of the quark and gluon momentum distributions in a fast moving hadron, aka the quasi distributions [6]. As an important advantage, it bypasses the inverse problem (IP) of reconstructing the PDF $x$-dependence from correlations of limited range, which afflicts other theoretical approaches such as the short-distance factorization (SDF), including the pseudo-PDF [7, 173] and “good lattice cross section” [8]. The SDF approach has also gained notable traction and led to substantial progress in extracting the lowest Mellin moments and phenomenological fitting of collinear PDFs and GPDs [174, 173, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 85, 193, 194, 195, 196, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 110, 208, 209, 210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 160], despite challenges in controlling the model uncertainty in the latter and its limited applicability to a broader range of parton physics.
@@
Figure 1: An illustration of the difference between LaMET and SDF: LaMET in principle utilizes data at all $z$ to sharply resolve the momentum of a parton, though in practice they are usually truncated at $z_{L}\sim 1.0$ fm due to noisy data and extrapolated using physics-motivated asymptotic forms to $z=\infty$ [78]. Meanwhile, the SDF can only use data up to about $z_{\rm max}\sim 0.2-0.3$ fm (shaded blue region) to fit a model of the PDF, which constitutes an IP. Plotted are the hybrid-scheme nucleon transversity quasi-PDF matrix elements computed in Ref. [111]. The shaded-red ellipse approximately indicates the sub-asymptotic region where the exponential decay starts to dominate. The concern raised in [1] is whether the data precision in this region allows for a reliable error estimate of the extrapolation; otherwise, it must be treated as an IP.
@@
where $|P\rangle$ is a plane-wave hadron state with momentum $P^{\mu}=(P^{t},0,0,P^{z})$, and $W(z,0)$ with $z^{\mu}=(0,0,0,z)$ is a spatial Wilson line that guarantees the gauge invariance of the correlator. In SDF, the data contain leading-twist information up to about z=0.20.3z=0.2-0.3 fm, beyond which the framework breaks down. For instance, the average size of the instanton, a non-perturbative gluon configuration, is about 0.3 fm [231, 232]. At finite momentum $P^{z}$, only a limited range of leading-twist correlations can be accessed, which is insufficient for a reliable forward determination of the PDFs. Instead, this segment of correlations can be used for phenomenological fittings of the PDFs. In contrast, LaMET legitimately utilizes matrix elements at all $z$ to compute the quasi-PDF, thereby providing a sharply-defined momentum fraction $x$ for quarks and gluons. This explains why LaMET can calculate the PDFs at a fixed momentum fraction $x$ through a systematic expansion. This distinction between LaMET and SDF has been extensively discussed in the literature [12, 78, 233, 234, 172], where it is argued that, given the same lattice data at a large $P^{z}$, LaMET offers the most accurate determination of a PDF within a moderate range of $x$.
Recently, Ref. [1] raised practical concerns about whether lattice data in the sub-asymptotic region, i.e., $0.7\penalty 10000\ \text{fm}\lesssim z\lesssim 1.0\penalty 10000\ \text{fm}$ presented there, are precise enough for a controlled physical extrapolation to larger $z$. If not, this could adversely impact the Fourier transform (FT) used to obtain quasi-PDFs for the LaMET expansion, rendering the error estimation in the literature possibly unreliable. Various data-driven IP methods [235, 236, 237, 238, 239] were proposed to treat the FT, yielding large uncertainties of the reconstructed quasi-PDF, dubbed as “inverse problem in the LaMET framework”.
In this paper, we directly respond to these concerns. In Sec. II, we first briefly review the fundamental difference in the predictive capabilities of LaMET and SDF, based on the extensive discussions in Refs. [78, 12, 233, 234, 172], emphasizing LaMET as a forward-problem (FP) formalism of calculating, not fitting, parton distributions. Then in Sec. III we point out that the alleged IP in LaMET application to lattice data is a standard asymptotic extrapolation problem in lattice QCD, which can be systematically treated with reliable error estimates. In Sec. IV, we examine in detail the criticisms and data analysis in Ref. [1], and conclude that the observed FT errors can be bounded by a theoretical estimate based on asymptotic extrapolation, and the data-driven IP methods used are less reliable as they do not satisfy the physical constraints. We conclude in Sec. V. Finally, since the authors of Ref. [1] recently commented [240] on our current paper, we provide a response in App. A.
II LaMET As a Forward-Problem Formalism
@@
As presented in Eq. (3), the alleged IP in Ref. [1] is essentially a matter of convergence and precision of the numerical Fourier integral, which are influenced by the interpolation and extrapolation of lattice data points in $z$-space. The smoothness of correlation functions in a bounded hadron ensures that the interpolation errors are small, while their asymptotic behavior at large distances controls the extrapolation uncertainty.
Since perturbative matching is already encoded in an EFT, it does not introduce an additional IP in the region where the framework is valid. In the literature, the coefficient $C$ in Eq. (II) has been derived by inverting the kernel that matches the PDF onto the quasi-PDF[13, 45], which raised the question of whether this procedure constitutes an IP. To address this concern, we note that the leading-power matching kernel can be inverted order by order in $\alpha_{s}$, since the calculation is performed in the asymptotic limit $P^{z}\to\infty$ where perturbation theory is valid across the entire kinematic range. However, at finite $P^{z}$, both perturbation theory and the power expansion break down in certain kinematic regions. As a result, Eq. (II) and the expansion of the quasi-PDF in terms of the PDF correspond to two distinct EFT expansions, which cannot be transformed into one another. This comparison is analogous to that between the quasi- and pseudo-PDFs [78, 12, 233, 234, 172].
In practice, there may still be concern about whether the FT uncertainty at small-$x$ in the quasi-PDF will propagate to moderate $x$ after its convolution with the matching kernel in Eq. (II). First, matching is performed between the collinear modes in the quasi-PDF and those in the PDF [12, 233, 234, 172], with a hard momentum $xP^{z}$. In contrast, contributions from soft modes with $x\to 0$ are power suppressed.
This suppression can also be derived from the structure of the matching kernel $C(x/y)$. For example, up to next-to-leading order (NLO),
$C\Big({x\over y}\Big)$ $=\delta\Big({x\over y}-1\Big)+{\alpha_{s}\over 2\pi}C^{(1)}\Big({x\over y}\Big)\,$, (4)
$C^{(1)}\Big({x\over y}\Big)$ $=C^{(1)}_{r}\Big({x\over y}\Big)-\delta\Big({x\over y}-1\Big)\int_{-\infty}^{\infty}dy^{\prime}C^{(1)}_{r}(y^{\prime})\,.$ (5)
At leading order, matching does not change the quasi-PDF. At NLO, in the limit $|x/y|\gg 1$,
$\lim_{|x/y|\gg 1}C^{(1)}_{r}\Big({x\over y}\Big)\propto\Big({y\over x}\Big)^{2}\,.$ (6)
For finite $x$, the dominant contribution to the matching integral arises from the region $|y|\gtrsim|x|$, while contributions from $|y|\ll x$ are suppressed quadratically. In practice, the range of $x$ targeted in LaMET analyses lies well above the small-$x$ region where FT uncertainties become significant, rendering any contamination from the latter negligible.
Therefore, according to Eqs. (II) and (3), LaMET is an FP formalism in the lattice calculation of PDFs at moderate $x$, which is systematically improvable without an adverse IP.
It is important to emphasize that in this formulation, the shapes and mathematical properties of the PDFs are not assumed a priori, in contrast to the IP formulation for SDF [1]. In fact, PDFs are calculated one $x$ at a time, and the $x$-space functions come out as a prediction. Since no connections with partons are made before the large-momentum expansion, and the terms in SDF for the coordinate-space correlations—such as “higher-twist contamination at large zz[1]—are irrelevant. It is also unnecessary to invoke a notion of so-called smearing of order $O(\Lambda_{\rm QCD}/P_{z})$ in $x$-space [1] to understand how LaMET works; in the standard EFT framework, the relevant systematics are accurately described by the power accuracy, which begins at $O(\Lambda_{\rm QCD}^{2}/P_{z}^{2})$ in LaMET [241]. Moreover, any variation in the PDF—as hypothesized in Ref. [1]—must either be predicted by LaMET or lie within its overall uncertainty band. If an undetected large, sharp variation exists in the quasi-PDF, it would correspond to a persistent long-range mode in coordinate space, which is ruled out by color confinement. If such a feature were to appear only in the PDF at moderate $x$, it would require an equally large, oppositely signed variation in the power correction. This would contradict the convergence of the EFT expansion in the region where it is expected to hold. Importantly, this scenario can be readily tested by performing calculations at multiple values of $P^{z}$, yet no such behavior has been observed so far.
@@
Here $\cal C$} is the matching kernel, and the power corrections are suppressed by the short distance $z$. After taking into account higher-order perturbative corrections [96, 128, 241], the SDF is found to be reliable up to z=0.20.3z=0.2-0.3 fm, which corresponds to the strong coupling constant $\alpha_{s}(1/z)\sim 1$ and the breakdown of perturbation theory. As a result, the largest $\lambda$ that can be reached in contemporary lattice simulations is limited to $\lambda_{\rm max}\sim 3.0-4.0$, which is far from the asymptotic region where the light-cone correlation decays algebraically. Therefore, the $x$-dependence of PDFs cannot be calculated directly and must instead be fitted through parametrizations and Bayesian or neural-network approaches with assorted numerical and mathematical assumptions [173, 183, 189, 224].
@@
To facilitate the LaMET analysis, one needs to have coordinate-space correlation functions in a large-momentum hadron at all distances $z\in[0,\infty]$, as information at negative $z$ can be obtained from discrete symmetries. It is well known that lattice QCD cannot produce quality data for $z\gg 1$ fm due to the exponentially worsening signal-to-noise ratio (SNR). Therefore, we cannot have all the data that LaMET requires. Fortunately, physical principles offer important guidance. Even without doing a very challenging calculation, QCD predicts that the correlation function must effectively vanish at sufficiently large $z$ due to color confinement. In fact, more precise theoretical predictions are available. According to dispersion theory, the correlation function decays exponentially at large $z$ with the decay length controlled by the residual mass of a heavy-light meson [97]. Moreover, this mass is an intrinsic property of the correlation function in the sense that it is independent of the external hadron momentum. Through a more detailed analysis [242], one can derive subleading behaviors of the correlation function as well. Armed with these known rigid physics constraints, the lattice data can be reliably extrapolated to larger $z$ in the asymptotic region.
@@ III.1 Asymptotic Extrapolation
Asymptotic expansion is essentially a generalization of the Taylor expansion in small parameters, which is widely used in theoretical physics, such as EFTs. It is also one of the most frequently used techniques in lattice QCD analyses, where the physical conditions are imposed as a prior. The precision of calculations is often assessed based on this procedure.
@@
Other examples include continuum and chiral extrapolations [246, 247, 248], where subleading effects from finite lattice spacing and unphysical pion masses are derived using EFTs and incorporated into fits to lattice data. In the same spirit, since dispersion relation analysis is a rigorous and model-independent method in theoretical physics, it supports the use of asymptotic extrapolation for spatial correlation functions as a standard and justified procedure.